# Utilizing Large Language Models for Information Extraction from Real Estate Transactions

### v2

@@ Abstract
**footnotetext: These authors contributed equally to this work
1 Introduction
Real estate transactions involve complex legal documents, such as sales contracts, that outline terms and conditions agreed upon by buyers and sellers. Extracting key information from these contracts is essential for various purposes, including due diligence, risk assessment, and compliance. However, manual review and extraction of data from these documents can be labor-intensive and prone to errors.
Real estate transactions are uniquely complex due to several distinctive factors. One key aspect is the presence of contingencies, which are conditions that must be met for the transaction to proceed, such as financing or inspections. These contingencies add negotiation layers and uncertainty. Additionally, real estate transactions often involve an executory period spanning weeks or months, allowing time for inspections and repairs before final closing. Property ownership is transferred through a deed, a legal document that conveys ownership rights and must be carefully drafted and executed. Furthermore, transactions entail various liabilities like environmental issues or property defects, requiring disclosure and mitigation to minimize risk. These factors highlight the specialized expertise needed to navigate real estate transactions successfully.
In this paper, we investigate the use of large language models for extracting structured data from real estate sales contracts. Traditionally, LSTM recurrent networksSchmidhuber et al. (1997) and TransformersVaswani et al. (2017) have been widely utilized to analyze sequential data in a variety of domains. We explore techniques to preprocess documents, fine-tune models, and information extraction techniques.
2 Related Work
Recent advances in natural language processing (NLP) Chang et al. (2024), particularly with language models like BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2018) and GPT (Generative Pre-trained Transformer) Brown et al. (2020), offer promising solutions for automating the extraction of information from text. These models, which have been proven to excel in a number of applications requiring sophisticated context understanding and reasoningGao et al. (2024), excel in understanding and generating human-like text, making them suitable for complex document analysis tasks.
Traditional statistical and machine learning approaches such as Conditional Random Fields (CRF), Support Vector Machines (SVM) and Hidden Markov Model (HMM) have been explored to extract named entities from legal contracts Nadeau and Sekine (2007) Betts and Jaep (2016) Surden (2021) Cui et al. (2023) Zamani and Schwartz (2017). Joshi et al. applied a sequence of information retrieval and traditional machine learning methods to determine the governing law of a contract.
@@ 3 Motivations
Furthermore, LLMs offer valuable tools for real estate agents, buyers, and sellers to comprehend contract terms effectively. By translating legal jargon into layman’s terms, LLMs empower individuals without legal expertise to understand the key provisions and implications of the contract. This enhanced understanding fosters transparency and facilitates informed decision-making during the negotiation and execution of real estate transactions. Real estate agents can better advise their clients, and buyers/sellers can navigate contractual terms with confidence, leading to smoother and more successful transactions overall.
@@ 4 Methodology
4.1 Data and Feature Preprocessing
@@
Additionally, positional encodings Wu et al. (2021) Vaswani et al. (2017) can be incorporated to convey sequential information in the input. This involves computing sinusoidal positional encodings based on the position $pos$ and dimension $d$ of each token, defined as:
$\text{PE}(pos,2i)=\sin\left(\frac{pos}{k^{2i/d}}\right)$
$\text{PE}(pos,2i+1)=\cos\left(\frac{pos}{k^{2i/d}}\right)$
where $i$ denotes the dimension index within the embedding vector and $k$ denotes a tuning parameter controlling the sequential relationships between texts. It’s worth noting that real estate transactions exhibit less sequential interdependence, where each paragraph generally aligns with a specific clause in a contract or a particular point in a report, resulting in relatively independent paragraphs. These positional encodings complement token embeddings and aid in capturing sequential relationships within the contract text.
@@
4.2 Fine-tuning Large Language Models
Fine-tuning large language models (LLMs) for real estate contract text data involves various approaches that leverage transfer learning, task-specific fine-tuning, and multi-task learning to enhance model performance and enable domain-specific learning. Transfer learning involves using pre-trained LLMs, like BERTDevlin et al. (2018) or GPTAchiam et al. (2023), trained on massive datasets for general language understanding. These models are then fine-tuned on real estate contract text data to adapt their learned features and knowledge to the nuances of real estate transactions. The fine-tuning process updates model parameters to better align with the target domain, enhancing the LLM’s ability to understand real estate-specific language and concepts.
@@ 4.2 Fine-tuning Large Language Models
Multi-task learning is a technique that involves training an LLM on multiple related tasks simultaneously Mahabadi et al. (2021), including real estate-specific tasks like contract interpretation and clause extraction. This approach has been used in numerous applications in specific domains Chakrabarty et al. (2019) because it encourages the model to learn shared representations across tasks, enabling it to generalize better and improve performance on each individual taskHoward and Ruder (2018) Wallingford et al. (2022). Multi-task learning helps the LLM leverage common patterns and features across different real estate contract-related tasks, leading to enhanced domain-specific learning and more robust performance.
4.3 Information Extraction
@@
Mathematically, these approaches involve training LLMs with labeled real estate contract data $\mathcal{D}$ and learning to predict key information categories $\mathcal{C}$ based on the contract text $x$. Let $f(x;\theta)$ represent the LLM model parameterized by $\theta$, and $\hat{y}=f(x;\theta)$ be the predicted information categories. The objective is to minimize a loss function $\mathcal{L}(\hat{y},y)$ that measures the model’s performance in predicting the correct information categorie $y$ from the real estate contract text $x$.
5 Query
Once an LLM model is fine-tuned, it is capable of answering a wide range of questions related to real estate transactions. For example, queries like ”Describe the area of the property” can be effectively answered by the model, leveraging its understanding of geographic references and property descriptions. Similarly, questions about specific contract terms, such as ”What are the contingencies in the contract?” can be addressed based on the model’s training on legal language and contract structures. By fine-tuning the LLM with relevant real estate data and legal documents, the model gains the ability to interpret and respond to diverse inquiries related to property details, transaction terms, and legal provisions, providing valuable insights and information to users involved in real estate dealings. This versatility in answering questions contributes to the efficiency and effectiveness of utilizing LLMs in real estate transactions, enhancing accessibility and understanding across various aspects of property transactions and contracts.
In fact, real estate professionals can create a tailored set of questions encompassing the crucial due diligence questions for each transaction. These questions can subsequently be inputted into the model for every deal, enabling the extraction of pertinent information and presenting key details to stakeholders involved in the transaction.
6 Dataset Generation
@@
Synthetic data has emerged as a valuable tool for training Large Language Models (LLMs) due to its potential to overcome limitations associated with real-world data, such as privacy concerns, biases, and limited availability of specific examples. Several methods have been explored to leverage synthetic data for LLM trainingLiu et al. (2024):
@@
We obtained authentic real estate transaction datasets from public sources, covering real estate transactions in New York CityNYC (2017). Those datasets are in tabular formats, which can be easily processed and understood by machine learning algorithms, and the columns contain key attributes of real estate transactions, like city, lot number, transaction time, and prices.
@@
For clauses not available in the real transaction data, we adopted a rule engine as described in Figure LABEL:fig:_synthetic to generate random contract terms and use them as ground truth answers to evaluate the model’s reasoning capabilities.
@@
7 Model Training and Experiments
While the field of natural language processing has been dominated by massive language models with hundreds of billions of parameters, our work takes a different approach. We leverage a smaller large language model (LLM), with a scale of a few billion parameters, exemplified by models like LLaMA-8BTouvron et al. (2023) and Phi-3Abdin et al. (2024). This deliberate choice is motivated by several crucial factors that align with our application’s specific requirements and constraints. Firstly, deploying a smaller LLM significantly enhances privacy. Smaller models enable on-device processing, unlike their larger counterparts, which often necessitate cloud-based inference. This means sensitive user data remains on the user’s device, minimizing the risks associated with data transmission and storage on external servers. This is particularly critical in privacy-sensitive domains where data security is paramount. Secondly, smaller LLMs offer practical advantages in inference speed and fine-tuning. Their reduced scale translates to lower computational demands, making them easier to adapt and specialize using modest resources. This efficiency is particularly beneficial when working with synthetic datasets, as it allows for rapid experimentation and iterative refinement of the model to achieve optimal performance on the specific task.
To further enhance the efficiency and performance of our chosen LLM, we employ Parameter-Efficient Fine-Tuning (PEFT) techniquesMangrulkar et al. (2022). PEFT methods are designed to adapt large language models for specific tasks by fine-tuning only a small subset of parameters while keeping most of the model’s weights frozen. By limiting the number of trainable parameters, PEFT significantly reduces the computational resources required for fine-tuning. This makes it feasible to adapt large models even with limited hardware. PEFT methods minimize the memory footprint during fine-tuning, enabling the training process on devices with less memory capacity. It also preserves the majority of general knowledge existing in the pre-trained model parameters. Low-Rank Adaptation (LoRA)Hu et al. (2022) inserts adapters with trainable rank decomposition matrices into each layer of the transformer model. In our work, we carefully evaluate different PEFT methods to identify the most effective approach for fine-tuning our chosen LLM on the synthetic dataset. By combining the advantages of a smaller LLM with the efficiency of PEFT techniques, we balance performance, resource utilization, and privacy.
On top of Llama-3.1-8B Instruct modelDubey et al. (2024), we adopt the following LoRA hyper-parameter configs for fine-tuning, which results in a massive reduction of trainable parameters to only 10 million.
@@ 7 Model Training and Experiments
Figure 2: Loss curve
8 Evaluation
@@ 8.1 Evaluation Metrics
Although the F1 scores provide a quantitative measure of answer accuracy, they often fall short of capturing the semantic similarity between the predicted and ground truth answers. To address this limitation, we incorporate BERTScoreZhang* et al. (2020) as an evaluation metric for our QA system. Unlike string-based comparisons, BERTScore leverages pre-trained contextual embeddings from BERT to compute the similarity between two sentences. This allows for a more nuanced assessment that considers the semantic meaning and contextual information embedded within the text, rather than relying solely on lexical overlap. By utilizing BERTScore, we aim to evaluate the quality of generated answers based on their semantic alignment with the ground truth, even if they exhibit variations in phrasing or lexical choices. This approach provides a more comprehensive and meaningful evaluation of the LLM’s ability to understand and respond to questions accurately and comprehensively.
@@
Table 1: Evaluation metrics show significant improvements after fine-tuning LLM with synthetic dataset
8.2 Qualitative Analysis

### v3

@@ Abstract
Index Terms:
Artificial intelligence, machine learning, large language models
I Introduction
Real estate transactions involve complex legal documents, such as sales contracts, that outline the terms and conditions agreed upon by buyers and sellers. Extracting key information from these contracts is essential for various purposes, including due diligence, risk assessment, and compliance. However, manual review and extraction of data from these documents can be labor intensive and prone to errors.
Real estate transactions are uniquely complex due to several distinctive factors. One key aspect is the presence of contingencies, which are conditions that must be met for the transaction to proceed, such as financing or inspection contingencies. These contingencies add layers of negotiation and uncertainty. In addition, real estate transactions often involve an executory period that spans weeks or months, allowing time for inspections and repairs before final closing. Property ownership is transferred through a deed, a legal document that conveys ownership rights, and must be carefully drafted and executed. Furthermore, transactions entail various liabilities, such as environmental issues or property defects, requiring disclosure and mitigation to minimize risk. These factors highlight the specialized expertise needed to navigate real estate transactions successfully.
In this paper, we investigate the use of large language models for extracting structured data from real estate sales contracts. Traditionally, LSTM recurrent networks[1] and Transformers[2] have been widely used to analyze sequential data in a variety of domains. We explore techniques to preprocess documents, fine-tune models, and information extraction techniques.
II Related Work
Recent advances in natural language processing (NLP) [3], particularly with language models like BERT (Bidirectional Encoder Representations from Transformers) [4] and GPT (Generative Pre-trained Transformer) [5], offer promising solutions for automating the extraction of information from text. These models, which have been proven to excel in a number of applications requiring sophisticated context understanding and reasoning[6], excel in understanding and generating human-like text, making them suitable for complex document analysis tasks.
Traditional statistical and machine learning approaches such as Conditional Random Fields (CRF), Support Vector Machines (SVM) and Hidden Markov Model (HMM) have been explored to extract named entities from legal contracts [7, 8, 9, 10, 11]. [12] applied a sequence of information retrieval and traditional machine learning methods to determine the governing law of a contract.
@@ IV Methodology
IV-A Data and Feature Preprocessing
@@
To convey sequential information in the input, Rotary Position Embedding (ROPE) [13] can be incorporated. ROPE is a relative positional encoding technique that applies a rotation operation to the token embeddings, enabling the model to capture positional relationships directly in the attention mechanism. For a given token embedding $\mathbf{x}\in\mathbb{R}^{d}$ at position pospos, ROPE applies a rotation matrix $\mathbf{R}(pos)$, transforming the embedding as follows:
$\text{ROPE}(\mathbf{x},pos)=\mathbf{R}(pos)\cdot\mathbf{x},$
where $\mathbf{R}(pos)$ is a block-diagonal matrix that rotates embedding dimensions in pairs. This approach integrates relative positional information into the model without introducing explicit positional vectors into the embedding layer.
Real estate transactions often exhibit less sequential interdependence, as each paragraph in a contract or report typically corresponds to a distinct clause or specific point. ROPE is well-suited for such data, as it efficiently encodes positional relationships while preserving flexibility for processing relatively independent sections of text. By encoding positional relationships implicitly, ROPE enhances the model’s ability to understand both local and global contexts in real estate contracts.
@@
IV-B Fine-tuning Large Language Models
Fine-tuning large language models (LLMs) for real estate contract text data involves various approaches that leverage transfer learning, task-specific fine-tuning, and multi-task learning to enhance model performance and enable domain-specific learning. Transfer learning involves using pre-trained LLMs, like BERT[4] or GPT[14], trained on massive datasets for general language understanding. These models are then fine-tuned on real estate contract text data to adapt their learned features and knowledge to the nuances of real estate transactions. The fine-tuning process updates model parameters to better align with the target domain, enhancing the LLM’s ability to understand real estate-specific language and concepts.
@@ IV-B Fine-tuning Large Language Models
Multi-task learning is a technique that involves training an LLM on multiple related tasks simultaneously [15], including real estate-specific tasks like contract interpretation and clause extraction. This approach has been used in numerous applications in specific domains [16] because it encourages the model to learn shared representations across tasks, enabling it to generalize better and improve performance on each individual task[17] [18]. Multi-task learning helps the LLM leverage common patterns and features across different real estate contract-related tasks, leading to enhanced domain-specific learning and more robust performance.
IV-C Information Extraction
@@
Mathematically, training large language models (LLMs) involves learning to predict the next token in a sequence of real estate contract text $x$. The dataset $\mathcal{D}$ consists of labeled real estate contract data, and the goal is to predict key information categories $\mathcal{C}$ token by token. Let $f(x;\theta)$ represent the LLM parameterized by $\theta$, and let the predicted probability distribution over the vocabulary for the next token $x_{t+1}$ be $\hat{y}_{t+1}=f(x_{t};\theta)$.
The training objective is to minimize the negative log-likelihood of the correct token $x_{t+1}$ under the predicted probability distribution:
$\mathcal{L}(\theta)=-\sum_{t}\log P(x_{t+1}\mid x_{\leq t};\theta),$
where $P(x_{t+1}\mid x_{\leq t};\theta)$ is obtained via a softmax over the model’s logits for the next token.
During generation, the model operates in an autoregressive pattern, where each predicted token $x_{t+1}$ is conditioned on all previous tokens $x_{\leq t}$. This process continues until an end-of-sequence token is generated, enabling the extraction of structured information from real estate contracts in a sequential manner.
V Query
Once an LLM model is fine-tuned, it is capable of answering a wide range of questions related to real estate transactions. For example, queries like ”Describe the area of the property” can be effectively answered by the model. Similarly, questions about specific contract terms, such as ”What are the contingencies in the contract?” can be addressed based on the model’s training on legal language and contract structures. By fine-tuning the LLM with relevant real estate data and legal documents, the model gains the ability to interpret and respond to diverse inquiries related to property details, transaction terms, and legal provisions, providing valuable insights and information to users involved in real estate transactions.
In fact, real estate professionals can create a customized set of questions that encompass key due diligence questions. These questions can be fed into the model for each transaction to automatically extract information for an arbitrary transaction.
VI Dataset Generation
@@
Synthetic data has emerged as a valuable tool for training Large Language Models (LLMs) due to its potential to overcome limitations associated with real-world data, such as privacy concerns, biases, and limited availability of specific examples. Several methods have been explored to leverage synthetic data for LLM training[19]:
@@ VII Experiment Setup
We obtained authentic real estate transaction datasets from public sources, covering real estate transactions in New York City[20]. Those datasets are in tabular formats, which can be easily processed and understood by machine learning algorithms, and the columns contain key attributes of real estate transactions, like city, lot number, transaction time, and prices.
@@
For clauses not available in the real transaction data, we adopted a rule engine as described in Figure 1 to generate random contract terms and use them as ground truth answers to evaluate the model’s reasoning capabilities.
@@
VIII Model Training and Experiments
While the field of natural language processing has been dominated by massive language models with hundreds of billions of parameters, our work takes a different approach. We leverage a smaller large language model (LLM), with a scale of a few billion parameters, exemplified by models like LLaMA-8B[21] and Phi-3[22]. This deliberate choice is motivated by several crucial factors that align with our application’s specific requirements and constraints. Firstly, deploying a smaller LLM significantly enhances privacy. Smaller models enable on-device processing, unlike their larger counterparts, which often necessitate cloud-based inference. This means sensitive user data remains on the user’s device, minimizing the risks associated with data transmission and storage on external servers. This is particularly critical in privacy-sensitive domains where data security is paramount. Secondly, smaller LLMs offer practical advantages in inference speed and fine-tuning. Their reduced scale translates to lower computational demands, making them easier to adapt and specialize using modest resources. This efficiency is particularly beneficial when working with synthetic datasets, as it allows for rapid experimentation and iterative refinement of the model to achieve optimal performance on the specific task.
To further enhance the efficiency and performance of our chosen LLM, we employ Parameter-Efficient Fine-Tuning (PEFT) techniques[23]. PEFT methods are designed to adapt large language models for specific tasks by fine-tuning only a small subset of parameters while keeping most of the model’s weights frozen. By limiting the number of trainable parameters, PEFT significantly reduces the computational resources required for fine-tuning. This makes it feasible to adapt large models even with limited hardware. PEFT methods minimize the memory footprint during fine-tuning, enabling the training process on devices with less memory capacity. It also preserves the majority of general knowledge existing in the pre-trained model parameters. Low-Rank Adaptation (LoRA)[24] inserts adapters with trainable rank decomposition matrices into each layer of the transformer model. In our work, we carefully evaluate different PEFT methods to identify the most effective approach for fine-tuning our chosen LLM on the synthetic dataset. By combining the advantages of a smaller LLM with the efficiency of PEFT techniques, we balance performance, resource utilization, and privacy.
On top of Llama-3.1-8B Instruct model[25], we adopt the following LoRA hyper-parameter configs for fine-tuning, which results in a massive reduction of trainable parameters to only 10 million.
@@ VIII Model Training and Experiments
Figure 2: Fine-tuning can significantly reduce training loss and validation loss, and the validation loss converges at the end of epoch 3.
IX Evaluation
@@ IX-A Evaluation Metrics
Although the F1 scores provide a quantitative measure of answer accuracy, they often fall short of capturing the semantic similarity between the predicted and ground truth answers. To address this limitation, we incorporate BERTScore[26] as an evaluation metric for our QA system. Unlike string-based comparisons, BERTScore leverages pre-trained contextual embeddings from BERT to compute the similarity between two sentences. This allows for a more nuanced assessment that considers the semantic meaning and contextual information embedded within the text, rather than relying solely on lexical overlap. By utilizing BERTScore, we aim to evaluate the quality of generated answers based on their semantic alignment with the ground truth, even if they exhibit variations in phrasing or lexical choices. This approach provides a more comprehensive and meaningful evaluation of the LLM’s ability to understand and respond to questions accurately and comprehensively.
@@
TABLE I: Evaluation metrics show significant improvements after fine-tuning LLM with synthetic dataset
@@
IX-B Qualitative Analysis